The load L that a horizontal beam can safely support is directly proportional to the square of the depth, d, of the beam. A beam with a depth of 8 inches can safely support 800 pounds.
a)Find the constant variation and write the variation equation
b)Find the load that a similar beam with a depth of 6 inches can safely support.
I set this up at a cross multiply problem and got the wrong answer, so I'm not sure where I went wrong. As for b, I'm not sure how to even go about solving that. Thanks so much!

Answer by stanbon(75887)

The load L that a horizontal beam can safely support is directly proportional to the square of the depth, d, of the beam. A beam with a depth of 8 inches can safely support 800 pounds.
a)Find the constant variation and write the variation equation
L = k*d^2
800 = k*8^2
k = 800/64 = 12.5
----
That is the constant of variation.
=======================================
b)Find the load that a similar beam with a depth of 6 inches can safely support.
L = 12.5*d^2
L = 12.5*6^2
Load = 450 pounds
======================
cheers,
Stan H.
=====================